OpenStudy (neonumbrella5115):

Why is 0 to the power of 1 equal to 1?

OpenStudy (anonymous):

Do you mean anything raised to the power of 0? Like \[1^{0}\]?

OpenStudy (aum):

0 to the power of 1 = \(0^1 = 0\) 1 to the power of 0 = \(1^0 = 1\)

OpenStudy (superhelp101):

umm because if mult. 1 0 times you get1

OpenStudy (aum):

\[ \large 1 = \frac{a^n}{a^n} = a^{n-n} = a^0 \\ \large a^0 = 1; ~~~1^0 = 1, ~~~~2^0 = 1; ~~~~.... \]

OpenStudy (anonymous):

Everything raised to the power of 0 is zero

OpenStudy (neonumbrella5115):

@Concentrationalizing : Yes, but I also mean 0^1 @aum Thank you!